find the first the terms of each arithmetic series.
first number= 14 last number= -85 sum of first n numbers= 900

Answer by stanbon(75887)

find the first the terms of each arithmetic series.
first number= 14 last number= -85 sum of first n numbers= -900
--------------
S(n) = (n/2)(1st + last)
---
-900 = (n/2)(14 + (-85))
-1800 = n(-71)
n = -1800/-71
==================
n has to be a whole number, and this is not.
=================================================
Cheers,
Stan H.